OpenStudy (xmoses1):

Solve the linear system. x + y = -3 y = 2x (-1, -2) no solution coincident

OpenStudy (anonymous):

dude help meeeeee plz member me

OpenStudy (anonymous):

let x+y=-3 y=2x becomes substitution: x+(2x)=-3 solve for "x" (3x)=-3 (x)=-1 x=-1 thus y=2x x=-1 by substitution y=2(-1) y=-2 hence, the answer is..... (-1,-2)! Happy Studying!!!!!

OpenStudy (abb0t):

x + 2x = -3 solve for x and plug it into y=2\(\sf \color{red}{x}\)

OpenStudy (anonymous):

Do you understand how I came to that answer?????

OpenStudy (xmoses1):

Yeah i am walking through the steps that you wrote. It makes perfect sense :)

OpenStudy (anonymous):

@abbot came a bit too late sorry!!!!! See yah on the flip side!

OpenStudy (xmoses1):

Thanks guys :)